OpenStudy (anonymous):

Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. Answer form must be A=(XXX/X PI + XX SQRT X) inches sq

OpenStudy (anonymous):

can someone help me please? gold medle and fan

OpenStudy (anonymous):

Okay @aum, But what formula do i use to figure out the question

OpenStudy (aum):

Let me redraw the figure so it is clearer: 

OpenStudy (aum):

What is the angle made by the chord AB at the center? That is, angle AOB = ?

OpenStudy (anonymous):

8*8*8?

OpenStudy (aum):

No, I am asking for the ANGLE made by the chord AB at the center of the circle.

OpenStudy (aum):

If three sides of a triangle are all of equal lengths what is that triangle called?

OpenStudy (anonymous):

equallateral triangle. so each angle must be 60

OpenStudy (aum):

Correct. Now convert 60 degrees to radians.

OpenStudy (anonymous):

divide them in half? so 30?

OpenStudy (aum):

No. How do you convert an angle in degrees to radians?

OpenStudy (aum):

How many degrees in a full circle? How many radians in a full circle?

OpenStudy (anonymous):

360 in a full circle 2 ?

OpenStudy (aum):

360 degrees in a full circle \(2\pi\) radians in a full circle. So 360 degrees is \(2\pi\) radians or 180 degrees is \(\pi\) radians. To convert degrees to radians, multiply by \(\pi\) and divide by 180.

OpenStudy (anonymous):

would it be a decimal

OpenStudy (aum):

Sector AOB is the shaded area: